OpenStudy (anonymous):

The sum of two numbers is 50. One number is 10 less than three times the other number. What are the two numbers?

OpenStudy (***[isuru]***):

take the 2 numbers as "x" and "y" what is x + y = ?

OpenStudy (anonymous):

I could say 25 x 25 but I don't think that is 10 less than 3 times the other number...

OpenStudy (***[isuru]***):

also they say that one number (x) is 10 less than the three times the other (3y) which means x = 3y - 10 u can write this as 3y - x = 10-----(1) and the sum of 2 numbers r 50 so.. x + y = 50 -----(2) now can u solve this system of equations ?

OpenStudy (anonymous):

let first number be "x" other number will be "3x-10" total of the two is 50 so, x+(3x-10)=50 4x-10=50 4x=60 x=15 the first number second number = 3x-10= 35 is the second number so the numbers are 15 & 35